OpenStudy (anonymous):

Please Check! Medal and Fan Papa's Italian Restaurant has collected data about customer sauce orders. It calculated that P(marinara) = 0.74, P(pesto) = 0.62, and P(marinara or pesto) = 0.90. Determine the P(marinara and pesto). A. 0.12 B. 0.46 C. 0.16 D. 0.22

OpenStudy (anonymous):

I think it is B.

hartnn (hartnn):

B is correct! :)

hartnn (hartnn):

\(P(A \cap B) = P(A)+P(B)- P(A \cup B)\)

OpenStudy (anonymous):

yay! :D thanks! :D My work: 0.9 = 0.74 + 0.62 - P(marinara and pesto) 0.9 = 1.36 - P(marinara and pesto) P(marinara and pesto) = 1.36 - 0.9 = 0.46 Does that looks right?

hartnn (hartnn):

absolutely :)

OpenStudy (anonymous):

thanks so much! :D

hartnn (hartnn):

welcome ^_^